I can go back and graph what i see out of problems but not write the equation correctly. how do i write an equation for a hyperbola in standard form with vertices (0,8) (0,-8) and asymptotes y=positive or negative 4x?
Answer by stanbon(48581)

write an equation for a hyperbola in standard form with vertices (0,8) (0,-8) and asymptotes y=positive or negative 4x
------------
Plot the two points and graph the two lines on a coordinate system.
--------------------------------
You should see that the center is (0,0)
Distance from center to vertex is a=8
Slope of the asymptote is a/b = 8/b = 4
That implies that b = 2.
--------------------------
EQUATION:
-(x^2)/4 + (y^2)/64 = 1
=================
Cheers,
Stan H.